### v2

# A priori estimates for parabolic Monge-Ampère type equations

Abstract.
We prove the existence and regularity of convex solutions to the first initial-boundary value problem for the parabolic Monge-Ampère equation
{$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $\psi,\phi$ are given functions, $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega\subset\mathbb{R}^{n}$ is a uniformly convex domain. Our approach can also be used to prove similar results for the $\gamma$-Gauss curvature flow with any $0<\gamma\leq 1$.
Key words and phrases:
Parabolic Monge-Ampère equation, Gauss Curvature Flow, regularity.
2010 Mathematics Subject Classification:
35K96, 35B65.
1. Introduction
The Monge-Ampère equation arises in various geometric problems, including the Minkowski problem and geometric flows. The elliptic Monge-Ampère equations have been extensively studied since the 1970s; we refer readers to [7, 25] and references therein for further details.
Over the past decades, the following parabolic Monge-Ampère equation has been widely investigated:
(1.1) {$-u_{t}+\rho(\det D^{2}u)&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a convex domain in $\mathbb{R}^{n}$.
To our best knowledge, the literature on this equation can be categorized into three main cases:
  • $(i)$
    $\rho(s)=s^{\frac{1}{n}}$.
  • $(ii)$
    $\rho(s)=-1/s$ and $\psi\equiv 0$;
  • $(iii)$
    $\rho(s)=\log s$.
In case $(i)$, Ivochkina and Ladyzhenskaia [14] obtained the a priori $\tilde{C}^{2,\alpha}$ estimates of the first initial-boundary value problem (see also [20]) if $\Omega$ and $\phi(\cdot,0)$ are uniformly convex, and if either
(1.2) $\inf_{\partial_{p}Q_{T}}\phi_{t}+\inf_{Q_{T}}\psi>\frac{1}{2}ad^{2}$
or
[…]
where $a=\max\{\sup_{Q_{T}}\psi_{t},0\}$ and $d=\text{diam}\,\Omega$. The a priori $C^{1,\alpha}$ and $W^{2,p}$ estimates for a generalized form of this equation were established in [13, 4]. Boundary regularity results were subsequently obtained by Tang [23]. Caffarelli-Schauder type estimates were derived in [29] under the assumption that $\psi$ is Lipschitz continuous.
For case $(ii)$, the equation was used by Tso [26] to investigate the Gauss curvature flow, where he derived $\tilde{C}^{2,\alpha}$ a priori estimates. The Schauder estimates and $W^{2,p}$ estimates were later established in [28] and [12], respectively. The Jörgens-Calabi-Pogorelov type results on entire solutions were obtained in [11, 31]. Moreover, one can refer to [20, 16] for another type of parabolic Monge-Ampère equation $-u_{t}\det D^{2}u=\Psi$.
In case $(iii)$, Tso [27] studied the first initial-boundary value problem and derived the $\tilde{C}^{2,\alpha}$ a priori estimates. Xiong and Bao [30] considered a generalized form where $\rho(s)=\eta(\log s)$ satisfies $\eta^{\prime}>0$ and $n\eta^{\prime\prime}\leq\eta^{\prime}$ to make the parabolic operator
$P(-u_{t},D^{2}u):=-u_{t}+\rho(\det D^{2}u)$
concave with respect to $D^{2}u$. Schnürer and Smoczyk investigated the Neumann boundary problem for a similar equation in [22].
However, the seemingly simplest scenario $\rho(s)=s$ remains less understood. Unlike the three cases discussed above, the parabolic operator $P(-u_{t},D^{2}u)=-u_{t}+\det D^{2}u$ is not concave. The existence of a global smooth solution of (1.1) in this case is a long-standing open problem due to the lack of concavity.
In a celebrated work [1], Caffarelli, Nirenberg, and Spruck established global $C^{2,\alpha}$ estimates for the Dirichlet problem of elliptic Monge-Ampère equations $\det D^{2}u=\psi$, proving the existence of unique smooth convex solutions. This naturally raises the question of whether the first initial-boundary value problem of parabolic Monge-Ampère equation
(1.4) {$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
has such global regularity under suitable conditions. In this article, we provide an affirmative answer to this question under conditions (1.3) given by Ivochkina and Ladyzhenskaia [14]. Specifically, we assume that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary, and $\psi,\phi$ are smooth functions satisfying the following conditions:
  • $(\mathcal{P}1)$
    $\phi_{t}+\psi\geq c_{0}>0$ on $\partial\Omega\times[0,T]$;
  • $(\mathcal{P}2)$
    $\phi(x,0)$ is uniformly convex with respect to $x$;
  • $(\mathcal{P}3)$
    $\psi(x,t)$ is concave with respect to $x$.
In this paper, we consider solutions in the following Banach space.
Definition 1.1.
For $Q=\Omega\times I$, where $\Omega\subset\mathbb{R}^{n}$ is an open convex domain, and $I=(a,b]$, we denote by $\tilde{C}^{k,\alpha}(\overline{Q})$ the Banach space with the norm
$||w||_{\tilde{C}^{k,\alpha}(\overline{Q})}=\sup_{\begin{subarray}{c}|\gamma|+2s\leqq k\\ (x,t)\in Q\end{subarray}}\left|D_{x}^{\gamma}D_{t}^{s}w(x,t)\right|+\sup_{\begin{subarray}{c}|\gamma|+2s=k\\ (x,t),(y,\tau)\in Q\end{subarray}}\frac{\left|D_{x}^{\gamma}D_{t}^{s}w(x,t)-D_{y}^{\gamma}D_{\tau}^{s}w(y,\tau)\right|}{\left(|x-y|^{2}+|t-\tau|\right)^{\alpha/2}}.$
We shall prove
Theorem 1.2.
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be convex in $x$ and solve equation (1.4) under the above assumptions. Then it holds
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{T})}\leq C,$
for some positive constants $C$ and $\alpha\in(0,1)$ depending only on $n,T,\Omega,c_{0},||\psi||_{\tilde{C}^{2}(\overline{Q}_{T})},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
By the a priori estimate, we obtain the existence of solution to equation (1.4).
Theorem 1.3.
Assume that equation (1.4) satisfies the compatibility condition of any order (see Remark 3.2) and the above assumptions. Then there exists a unique smooth solution $u$, convex in $x$, to the first initial-boundary value problem (1.4).
Remark 1.4.
It is worth pointing out that in subsection 3.2, we take examples to show that condition $({\mathcal{P}3})$ is necessary for the existence of convex solutions to problem (1.4).
We can also prove Theorems 1.2 and 1.3 for Gauss curvature flow with boundary
(1.5) {$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{(n+1)/2}}\quad\quad\text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\quad\quad\;\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
Equation (1.5), corresponding to this setting, was studied in [21], where the author proved the existence of certain self-similar solutions and obtained gradient estimates in the homogeneous case $\phi\equiv 0$. However, higher-order global regularity for this case remains an open problem.
Our main difficulty in proving Theorem 1.2 lies in proving the interior second-order derivative estimates and the boundary $\tilde{C}^{2,\alpha}$ estimates under the lack of concavity. It is worth pointing out that our approach to obtaining such global estimates requires a combination of corresponding boundary estimates for equation (1.4) and interior estimates for equation (1.8), the equation of the Legendre transform of $u$, rather than studying only one of the equations to obtain them.
The proof of Theorem 1.2 can be divided into the following three steps:
Step 1:  Establish the $C^{2}$ estimate on the boundary
We first obtain the $C^{0}$ and $C^{1}$ estimates, and
(1.6) $C^{-1}\leq u_{t}+\psi\leq C,$
for some positive constant $C$. Then, we obtain the bound of $D^{2}u$ on the parabolic boundary by using the method in [1]. Thus, by (1.6) we have the estimates for the eigenvalues
(1.7) $\lambda(D^{2}u)\geq C\quad\text{on}\quad\partial_{p}Q_{T}.$
Step 2:  Establish the global $C^{2}$ estimate by the duality method
Due to the lack of concavity, we cannot use the maximum principle to control the global maximum of $|D^{2}u|$ by $\sup_{\partial_{p}Q_{T}}|D^{2}u|$. Instead, we apply the estimates in Step 1 to the equation after the Legendre transform.
Let $U$ be the Legendre transform of $u$, given by
$U(y,t)=\sup\{y\cdot x-u(x,t):\ x\in\Omega\}.$
Denote $Q_{T}^{*}=\{(y,t):\,y\in Du(\Omega,t),t\in(0,T]\}$. Then $U$ satisfies
(1.8) $-U_{t}-\frac{1}{\det D^{2}U}=-\psi(DU,t)\quad\text{ in }Q_{T}^{*}.$
Since the above equation is concave, one can infer
$\sup_{Q_{T}^{*}}||D^{2}U||\leq\sup_{\partial_{p}Q_{T}^{*}}||D^{2}U||,$
which combined with (1.6) and (1.7) yields an upper bound for the eigenvalues $\lambda(D^{2}U)$ in $Q^{*}_{T}.$ Then, by duality we have $\lambda(D^{2}u)\geq C~{}\text{in}~{}Q_{T}.$ By (1.6), it follows that
(1.9) $C^{-1}\leq\lambda(D^{2}u)\leq C\quad\text{in}\quad Q_{T}.$
Step 3:  Establish the global $\tilde{C}^{2,\alpha}$ estimate
According to Step 2, equation (1.8) is uniformly parabolic. Since the equation is also concave, we can obtain the space Hölder estimate for $D^{2}U(\cdot,t)$ for each $t$ by the Evans-Krylov theorem. The interior $\tilde{C}^{2,\alpha}$ regularity then follows by [6, 17] and the interior higher-order regularity of $U$ also follows by the bootstrap technique.
As for the boundary estimate, we first use Krylov-Safonov’s estimate (see [18]) to get the global space-time Hölder estimate for $u_{t}$. Then by differentiating the equations along tangential directions $\tau_{k}$, we shall prove the $\tilde{C}^{1,\alpha}$ regularity of $D_{\tau_{k}}u$ on the lateral surface of the parabolic boundary, followed by the space-time Hölder estimate for second normal derivatives.
This paper is arranged as follows. In Section 2, we establish the a priori estimates in $\tilde{C}^{2,\alpha}(\Omega\times[\epsilon,T])$. We then obtain the existence and uniqueness of solutions in Section 3, completing the proof of Theorems 1.2 and 1.3. Finally, we prove the existence of a unique smooth solution to equation (1.5) in Section 4.
2. The a priori estimate
We first consider the parabolic Monge-Ampère equation
(2.1) {$-u_{t}+\det D^{2}u&=\psi(x,t)\quad\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\,\ \text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $Q_{T}=\Omega\times(0,T]$, $\partial_{p}Q_{T}$ is the parabolic boundary of $Q_{T}$, and $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\psi,\phi$ are smooth functions satisfying assumptions $(\mathcal{P}1)-(\mathcal{P}3)$.
In this section, we suppose that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is convex in $x$ and a solution to equation (2.1). We aim to prove the $\tilde{C}^{2,\alpha}$ a priori estimate for $u$. For this purpose, we shall obtain the $C^{2}$ estimate for $u$ on the parabolic boundary first.
Lemma 2.1 (Comparison Principle).
Assume that $w,v\in\tilde{C}^{2}(\overline{Q}_{T}\setminus\partial_{p}Q_{T})\cap C(\overline{Q}_{T})$ and one of $w,v$ is convex in $x$. If $-w_{t}+\det D^{2}w\geq-v_{t}+\det D^{2}v$ in $\overline{Q}_{T}\setminus\partial_{p}Q_{T}$ and $w\leq v$ on $\partial_{p}Q_{T}$, then $w\leq v$ in $Q_{T}$.
Proof.
See [20, Theorem 14.1]. ∎
2.1. Step 1: $C^{2}$ estimates on the boundary
At first, we obtain the $C^{0}$ and $C^{1}$ estimates, stated as the lemma below.
Lemma 2.2.
$u(\cdot,t)$ is strictly convex for any $t\in[0,T]$, and the following estimates hold,
(2.2) $||u||_{C^{1}(\overline{Q}_{T})}\leq C,$
and
(2.3) $1/C\leq u_{t}+\psi\leq C.$
Here $C$ is a universal positive constant that depends only on $n,T$, $\Omega$, $c_{0}$, $||\phi||_{\tilde{C}^{2}(\overline{Q}_{T})}$, $\|\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi_{t}\|_{L^{n+1}({Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
Proof.
By the comparison principle, we have
±$\pm u\leq\Big{(}\sup_{\partial_{p}Q_{T}}|\phi|+\sup_{{Q}_{T}}|\psi|\Big{)}e^{t}\quad\text{in}\ Q_{T},$
which yields that $\|u\|_{C^{0}(\overline{Q}_{T})}\leq(\sup_{\partial_{p}Q_{T}}|\phi|+\sup_{{Q}_{T}}|\psi|)e^{T}$.
To obtain (2.3), differentiating the equation (2.1) with respect to $t$, we have
(2.4) $\mathcal{L}u_{t}=\psi_{t}\quad\text{ in }\ Q_{T},$
where the linearized operator $\mathcal{L}$ of (2.1) is
(2.5) $\mathcal{L}=-\partial_{t}+\det D^{2}u\cdot u^{ij}\partial_{ij},$
and $u^{ij}$ is the $(i,j)$-entry of the inverse matrix of $D^{2}u$. Since $\psi(x,t)$ is concave in $x$, then
$\mathcal{L}(u_{t}+\psi)=\det D^{2}u\cdot u^{ij}\partial_{ij}\psi\leq 0.$
It follows form the weak maximum principle that
$\inf_{Q_{T}}(u_{t}+\psi)=\inf_{\partial_{p}Q_{T}}(u_{t}+\psi)=\min\{\inf_{\partial\Omega\times(0,T]}(\phi_{t}+\psi),\inf_{\Omega\times\{0\}}(\det D^{2}\phi)\}.$
That is
(2.6) $\inf_{Q_{T}}(u_{t}+\psi)\geq\min\{c_{0},\inf_{\Omega}(\det D^{2}\phi(\cdot,0))\}>0,$
or,
$\inf_{Q_{T}}u_{t}\geq-\sup_{Q_{T}}|\psi|+\min\{c_{0},\inf_{\Omega}(\det D^{2}\phi(\cdot,0)\}.$
By using the parabolic Alexandrov-Bakelman-Pucci maximum principle ([20, Theorem 7.1]) to equation (2.4), we have
[…]
where
$\Gamma_{u_{t}}=\{(x,t)\in Q_{T}:u_{tt}\geq 0,D^{2}u_{t}\leq 0\},\quad C_{n}=(n(n+1)/\omega_{n})^{\frac{1}{n+1}},$
and $\omega_{n}$ is the surface area of unit sphere in $\mathbb{R}^{n}$. Note that on $\Gamma_{u_{t}}$,
$\begin{split}u_{tt}\det(-D^{2}u_{t})&=(\det D^{2}u)^{-\frac{n-1}{n+1}}u_{tt}\cdot\det\Big{[}(\det D^{2}u)^{\frac{2}{n+1}}(D^{2}u)^{-1}(-D^{2}u_{t})\Big{]}\\ &\leq\left(\frac{u_{tt}+(\det D^{2}u)\text{tr}\Big{[}(D^{2}u)^{-1}(-D^{2}u_{t})\Big{]}}{(n+1)(\det D^{2}u)^{\frac{n-1}{n+1}}}\right)^{n+1}\\ &=\left(\frac{-\mathcal{L}u_{t}}{(n+1)(u_{t}+\psi)^{\frac{n-1}{n+1}}}\right)^{n+1}.\end{split}$
It follows that
[…]
Therefore, we have proved (2.3) according to (2.6) and (2.8). The strict convexity follows directly from (2.3) and the smoothness of $\phi$, see [2] for a reference.
To estimate $|Du|$, it suffices to prove the bound for $|D_{\nu}u|$ on $\partial\Omega\times[0,T]$ for the outward unit normal vector $\nu$, since $|Du|$ is bounded by $|D\phi|$ on $\Omega\times\{0\}$ and tangentially on $\partial\Omega\times[0,T]$. For any fixed $t_{0}\in[0,T]$, take a smooth auxiliary function $w(x)$ such that $w=\phi(t_{0})$ on $\partial\Omega$, and $\det D^{2}w\geq\sup_{Q_{T}}(u_{t}+\psi)$ in $\Omega$. Note that such $w$ does exist, and one example is to assume $w$ with $(-d_{\partial\Omega}+d^{2}_{\partial\Omega})\delta+\phi(t_{0})$ near the boundary for some constant $\delta>0$ large enough. Then by the classical comparison principle for Monge-Ampère equations, we have $u\geq w$ in $\Omega\times\{t_{0}\}$. Thus $D_{\nu}u\leq D_{\nu}w\leq C$. Till now we have proved formula (2.2). ∎
The proof for the second-order derivative estimates on the parabolic boundary follows a similar structure as in Section 2 of [1]. We first establish the boundedness of the second tangential derivatives of $u$ on the boundary, followed by the mixed second derivatives. The boundedness of the second normal derivatives is then derived from these estimates and the equation of $u$.
Theorem 2.3.
There is a positive constant $C$ such that
(2.9) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }\partial_{p}Q_{T},$
where $\lambda(D^{2}u)$ denotes the eigenvalues of $D^{2}u$ and the constant $C$ depends only on on $n,T$, $\Omega$, $c_{0}$, $||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$, $\|D\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi\|_{L^{\infty}({Q}_{T})}$, $\|\psi_{t}\|_{L^{n+1}({Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
Proof.
By the boundary condition, the matrix norm of $D^{2}u$ is bounded by $D^{2}\phi(\cdot,0)$ on $\Omega\times\{0\}$. We also note that by the uniform convexity of $\phi$, there exists a positive constant $C$ such that
$C^{-1}I_{n}\leq D^{2}u(x,0)\leq CI_{n}$
for $x\in\Omega$. Here $C$ depends on $||D^{2}\phi(\cdot,0)||_{L^{\infty}(\Omega)}$ and the uniform convexity of $\phi(\cdot,0)$. Let $(x_{0},t_{0})$ be an arbitrary point on $\partial\Omega\times[0,T]$, w.l.o.g. $x_{0}=0$ and $\partial\Omega$ be parameterized by $x_{n}=\rho(x^{\prime})$ on a small neighborhood of 00 with $\rho(0)=0$, $D\rho(0)=0$ and $\Omega\subset\{x_{n}>0\}$. We claim that
(2.10) $C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u(0,t_{0})\leq CI_{n-1}\quad\text{ for any }t_{0}\in[0,T].$
Differentiate the identity $u(x^{\prime},\rho(x^{\prime}),t)=\phi(x^{\prime},\rho(x^{\prime}),t)$ to obtain
(2.11) $(\nabla_{x^{\prime}}+\nabla\rho\partial_{n})(u-\phi)=0.$
Twice differentiation at point $(0,t_{0})$ gives that
$D_{x^{\prime}x^{\prime}}u(0,t_{0})=-u_{n}(0,t_{0})D^{2}\rho(0)+D_{x^{\prime}x^{\prime}}\phi(0,t_{0})+D^{2}\rho(0)\phi_{n}(0,t_{0}).$
We can obtain the latter inequality in (2.10) from this formula, as $|Du|,|D\phi|$ and $|D^{2}\phi|$ are bounded. The former inequality of (2.10) follows the same proof as in [1] if we regard $u_{t}+\psi$ as a data function at the time $t_{0}$, and the estimation depends on the maximum and minimum values of $u_{t}+\psi$, $\Omega$ and $\|\phi\|_{\tilde{C}^{4}{(\overline{Q}_{T})}}$.
Then we shall prove a uniform bound for the mixed second derivatives of $u$. Suppose that
(2.12) $\rho(x^{\prime})=\frac{1}{2}\sum_{\alpha,\beta<n}B_{\alpha\beta}x_{\alpha}x_{\beta}+O(|x^{\prime}|^{3})$
in a neighborhood of 00. Note that the matrix $B_{\alpha\beta}$ is positive definite because of the uniform convexity of $\partial\Omega$. Denote the operator
(2.13) $\mathcal{T}=\partial_{\alpha}+\sum_{\beta<n}B_{\alpha\beta}(x_{\beta}\partial_{n}-x_{n}\partial_{\beta}).$
Differentiating equation (2.1), we obtain
$\mathcal{L}\mathcal{T}u=\mathcal{T}\psi$
for the linear operator $\mathcal{L}$ defined in (2.5). Thus, by (2.3), we have
(2.14) $\begin{split}|\mathcal{L}\mathcal{T}(u-\phi)|&\leq|\mathcal{T}\psi|+|\mathcal{L}\mathcal{T}\phi|\\ &\leq C\|D_{x}\psi\|_{L^{\infty}(Q_{T})}+|(u_{t}+\psi)u^{ij}\partial_{ij}\mathcal{T}\phi-\partial_{t}\mathcal{T}\phi|\\ &\leq\underline{c}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}\end{split}$
for some positive constant $\underline{c}$ depending on $\Omega$, $\|D_{x}\psi\|_{L^{\infty}(Q_{T})}$, $||\phi||_{\tilde{C}^{3}(\overline{Q}_{T})}$ and the positive upper bound for $u_{t}+\psi$. Moreover, on $\Omega\times\{0\}$, $\mathcal{T}(u-\phi)=0$ and
(2.15) $|\mathcal{T}(u-\phi)|\leq C|x^{\prime}|^{2}|u_{n}-\phi_{n}|+\Big{|}\sum_{\beta<n}B_{\alpha\beta}x_{n}(u_{\beta}-\phi_{\beta})\Big{|}\leq\overline{c}x_{n},\quad\text{on }\partial\Omega\times[0,T],$
by the uniform convexity of $\Omega$ and the boundedness of $|Du|,|D\phi|$.
Consider the following barrier function
$w(x,t)=-a|x|^{2}+bx_{n}.$
Then $\mathcal{L}w=-2a(u_{t}+\psi)\sum_{i=1}^{n}u^{ii}$. Note that
$\sum_{i=1}^{n}u^{ii}\geq n\Big{(}\prod_{i=1}^{n}u^{ii}\Big{)}^{1/n}\geq n(\det D^{2}u)^{-1/n}\geq nC^{-1/n},$
where $C$ depends on the maximum of $u_{t}+\psi$. If we take $a$ large enough, then
$\mathcal{L}(w\pm\mathcal{T}(u-\phi))\leq 0.$
Since $\Omega$ is uniformly convex and bounded, there exists some universal constant $c$ such that $x_{n}\geq c|x|^{2}$ on $\overline{\Omega}$. Thus we could take $b\geq a/c+\overline{c}$ to derive
$w\pm\mathcal{T}(u-\phi)\geq 0\text{ on }\partial_{p}Q_{T}.$
By using the comparison principle for $u$ and $w$, we obtain that
$-w\leq\mathcal{T}(u-\phi)\leq w\quad\text{ in }\overline{Q}_{T}.$
Since $w(0,t)=\mathcal{T}(u-\phi)(0,t)=0$, this implies that
$|\partial_{n}\mathcal{T}(u-\phi)(0,t)|=|u_{\alpha n}(0,t)-\phi_{\alpha n}(0,t)|\leq b,$
which yields that
(2.16) $|u_{\alpha n}(0,t)|\leq b+||\phi||_{C^{2}(\overline{Q}_{T})}\leq C.$
Finally, we turn to the second normal derivatives of $u$. From the equation (2.1), we know that
$u_{t}(0,t)+\psi(0,t)=\det D_{x^{\prime}x^{\prime}}u(0,t)u_{nn}(0,t)+P\big{(}D_{x^{\prime}x^{\prime}}u(0,t),D_{x^{\prime}x_{n}}u(0,t)\big{)},$
where $P$ is a $n$-degree polynomial on $D_{x^{\prime}x^{\prime}}u(0,t)$ and $D_{x^{\prime}x_{n}}u(0,t)$. Combining formula (2.3), (2.10) and (2.16), we obtain a uniform bound for $u_{nn}$ on $\partial\Omega\times[0,T]$.
To conclude, we have proved a uniform bound for the second derivatives of $u$ on $\partial_{p}Q_{T}$, i.e.
(2.17) $|D^{2}u|\leq C\quad\text{ on }~{}~{}\partial_{p}Q_{T}.$
Since $u_{t}+\psi$ has a positive lower bound, then so does the smallest eigenvalue of $D^{2}u$ on the parabolic boundary according to (2.17). Thus estimate (2.9) is proved. ∎
2.2. Step 2: global $C^{2}$ estimates
Let $U$ be the Legendre transform of $u$ defined on $Q_{T}^{*}=\{(y,t):\,y\in Du(\Omega,t),t\in(0,T]\}$, i.e.,
$U(y,t)=\sup_{x\in\Omega}(y\cdot x-u(x,t)).$
Then $U$ satisfies
(2.18) $-U_{t}-\frac{1}{\det D^{2}U}=-\psi(DU,t)\quad\text{ in }Q_{T}^{*}.$
Note that $Du(\Omega,t)$ is not necessarily convex. Here $Q_{T}^{*}$ is bounded and simply connected due to the smoothness of $u$ and formula (2.2). By estimates (2.3) and (2.9) above, we know that
(2.19) $C^{-1}\leq\det D^{2}U\leq C\quad\text{and}\quad||D^{2}U||_{L^{\infty}(\partial_{p}Q_{T}^{*})}\leq C.$
Lemma 2.4.
If $U\in\tilde{C}^{4}(\overline{Q_{T}^{*}})$ satisfies equation (2.18), it holds that
$\sup_{Q_{T}^{*}}||D^{2}U||\leq\sup_{\partial_{p}Q_{T}^{*}}||D^{2}U||.$
Proof.
Differentiating the equation (2.18), we have for $k,l=1,\cdots,n$,
$-U_{tk}+\frac{1}{\det D^{2}U}U^{ij}U_{ijk}=-\psi_{i}U_{ik}$
and
$\begin{split}-U_{tkl}+\frac{1}{\det D^{2}U}U^{ij}U_{ijkl}-\frac{\big{(}U^{ij}U_{ijk}\big{)}\big{(}U^{ij}U_{ijl}\big{)}}{\det D^{2}U}-\frac{U_{ijk}U_{pql}U^{ip}U^{qj}}{\det D^{2}U}=-\psi_{ij}U_{ik}U_{jl}-\psi_{i}U_{ikl}.\end{split}$
If we define the linear operator $\tilde{L}=-\partial_{t}+\frac{1}{\det D^{2}U}U^{ij}\partial_{ij}+\psi_{i}\partial_{i}$, then
$\tilde{L}U_{kk}$ $=\frac{1}{\det D^{2}U}\Big{(}U^{ij}U_{ijk}\Big{)}^{2}+\frac{1}{\det D^{2}U}U_{ijk}U_{pqk}U^{ip}U^{qj}-\psi_{ij}U_{ik}U_{jk}.$
Since $U$ is convex and $\psi$ is concave, it holds that $U_{ijk}U_{pqk}U^{ip}U^{qj}\geq 0$ and $-\psi_{ij}U_{ik}U_{jk}\geq 0$. Therefore,
(2.20) $\tilde{L}U_{kk}\geq 0.$
By the comparison principle ([20, Theorem 2.4]) for linear parabolic equations, we obtain that
$U_{kk}\leq\sup_{\partial_{p}Q_{T}^{*}}U_{kk},$
or,
$\sup_{Q_{T}^{*}}||D^{2}U||\leq\sup_{\partial_{p}Q_{T}^{*}}||D^{2}U||.$
By formula (2.19) and Lemma 2.4, the eigenvalues of $D^{2}u$ have a positive lower bound in $Q_{T}$. Combined with (2.3), they are also bounded from above. To be more explicit, we have proved the following theorem:
Theorem 2.5.
It holds that
(2.21) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{ on }~{}~{}\overline{Q}_{T},$
where $C$ is a positive constant depending on those universal quantities in Theorem 2.3.
2.3. Step 3: global $\tilde{C}^{2,\alpha}$ estimates
In this part, we first give the interior $\tilde{C}^{2,\alpha}$ regularity for $u$.
Theorem 2.6.
Recall that $u\in\tilde{C}^{4}(\overline{Q}_{T})$ is a solution to equation (2.1) and is convex in $x$. Then, for any $Q\subset\subset Q_{T}$, there exists $\alpha\in(0,1)$ such that
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q})}\leq C,$
where $\alpha,C$ are positive constants depending on those universal quantities in Theorem 2.3 and dist ($\text{dist\,}(Q,Q_{T})$.
Proof.
Recall that $U_{kk}$ satisfies (2.20) as $\psi(x,t)$ is concave with respect to $x$. By using the same method in [26, Theorem 4.1], it holds that $U\in\tilde{C}^{2,\alpha}(\Omega\times(0,T])$. Then by (2.21), we have the interior regularity for $u$ in $Q_{T}$. ∎
Recall that
$\mathcal{L}u_{t}=\psi_{t}\quad\text{ in }\ Q_{T},$
where the linear operator $\mathcal{L}=-\partial_{t}+\det D^{2}u\cdot u^{ij}\partial_{ij}$ is defined in (2.5). By estimate (2.21), the operator $\mathcal{L}$ is uniformly parabolic. Thanks to the Krylov-Safonov’s estimate for linear parabolic equations (see [8, Theorem 5.1] or [18]), we can obtain the global $\tilde{C}^{\alpha}$ estimate to $u_{t}$.
Theorem 2.7.
Under the same conditions as in Theorem 2.6, we have
$||u_{t}||_{\tilde{C}^{\alpha}(\overline{Q}_{T})}\leq C$
for some positive constants $C$ and $\alpha$ depending on the uniformly parabolic coefficients of $\mathcal{L}$ and $\|\psi_{t}\|_{L^{n+1}(Q_{T})}$.
Remark 2.8.
The interior $\tilde{C}^{2,\alpha}$ regularity of $u$ can also be obtained by using the same method in [3, Proposition 5.3], once we get the global $C^{2}$ estimates and the $\tilde{C}^{\alpha}$ estimate for $u_{t}$.
We split the proof of the global $\tilde{C}^{2,\alpha}$ estimate into two parts. A short-time existence theorem gives the global $\tilde{C}^{\alpha}$ estimate for $D^{2}u$ on $\Omega\times(0,2\epsilon]$. As for the estimate on the rest part of $Q_{T}$, considering the linear parabolic equation satisfied by tangential derivatives of $u$, we use the approach of Theorem 9.31 in [10] to deduce the Hölder regularity of the gradient, which is independent of the moduli of continuity of the principal coefficients.
Let $Y_{0}=(y_{0},t_{0})$ be any point on $\partial\Omega\times[\epsilon,T]$ with some constant $\epsilon>0$. Without loss of generality, assume that $y_{0}=0$, and on a small neighborhood of 00, $\partial\Omega$ is the image of $y_{n}=\rho(y^{\prime})$ with $\rho(0)=0$ and $D\rho(0)=0$. Perform the transformation of coordinates $x=\Psi(y)$, and $\Psi:B_{\rho_{0}}(0)\cap\overline{\Omega}\to\overline{\mathbb{R}_{n}^{+}}$, with
$x^{\prime}=y^{\prime},\quad x_{n}=y_{n}-\rho(y^{\prime}).$
Let
$v(x,t)=T_{k}(u-\phi)(x^{\prime},x_{n}+\rho(x^{\prime}),t)$
for a fixed $1\leq k<n$, where
$T_{i}:=\partial_{y_{i}}+\rho_{y_{i}}\partial_{y_{n}}\quad\text{ for }~{}i<n,\quad T_{n}=\partial_{y_{n}}.$
Take $G=B_{R_{0}}^{+}\times(t_{0}-R_{0}^{2},t_{0}]$, such that $B_{R_{0}}^{+}\subset\operatorname{Im}\Psi$ and $R_{0}^{2}<\epsilon$. Then by direct computations,
$\mathcal{L}(T_{k}u)$ $=-\partial_{t}(T_{k}u)+\det D^{2}u\cdot u^{ij}(T_{k}u)_{ij}$
$=T_{k}\psi+\det D^{2}u\cdot(u^{ij}\rho_{kij}u_{n}+u^{ij}\rho_{ki}u_{nj}+u^{ij}\rho_{kj}u_{ni})=:\overline{f}.$
Denote $P_{i}=\partial_{x_{i}}-\rho_{x_{i}}\partial_{x_{n}}$ for $1\leq i<n$ and $P_{n}=\partial_{x_{n}}$. Thus under $(x,t)$ coordinates,
$-\partial_{t}v+\det D^{2}u\cdot u^{ij}P_{i}P_{j}v=Lv-\det D^{2}u\cdot u^{ij}\rho_{ij}v_{n}.$
Here the linear operator $L=-\partial_{t}+a^{ij}\partial_{x_{i}x_{j}}$ with
$a^{ij}=\det D^{2}u\cdot(u^{ij}+u^{sl}\rho_{s}\rho_{l}\delta_{n}^{i}\delta_{n}^{j}-u^{sj}\rho_{s}\delta_{n}^{i}-u^{is}\rho_{s}\delta_{n}^{j}),$
where $\delta_{n}^{i}$ equals $1$ when $n=i$, otherwise 00, and the summation over $s,l$ is from $1$ to $n-1$. Note that $P_{j}v=\partial_{y_{j}}(T_{k}(u-\phi))$ and $P_{i}P_{j}v=\partial_{y_{i}y_{j}}(T_{k}(u-\phi))$ for $i,j=1,\cdots,n$.
Let
$f(x,t)=\Big{(}\det D^{2}u\cdot u^{ij}\rho_{ij}\partial_{y_{n}}(T_{k}(u-\phi))-\mathcal{L}(T_{k}\phi)+\overline{f}\Big{)}(\Psi^{-1}(x),t).$
Then, $v$ satisfies the equation
(2.22) {$-v_{t}+a^{ij}v_{ij}&=f\quad\text{ in }G,\\ v&=0\quad\text{ on }F,\end{aligned}\right.$
where $F=\partial G\cap\{x_{n}=0\}$ is the flat boundary.
By Lemma 2.2 and Theorem 2.5, we have
(2.23) $\lambda I_{n}\leq(a^{ij})\leq\Lambda I_{n},$
if $\rho_{0}>0$ small, and
(2.24) $|f|\leq C_{0},\quad\quad|v|+|Dv|\leq C_{1}\quad\text{in}\quad\overline{G}$
for some positive constants $\lambda$, $\Lambda$, $C_{0}$ and $C_{1}$ depending on those universal quantities in Theorem 2.3.
Before proving the boundary $\tilde{C}^{1,\alpha}$ regularity of $v$, we first give three lemmas including the weak Harnack inequality, the interior $\tilde{C}^{1,\alpha}$ regularity of $v$, and a technical lemma.
Lemma 2.9.
Let $v\in\tilde{C}^{2}(\overline{G})$ be a solution of equation (2.22) with conditions (2.23) and (2.24). Then there are constants $10<\alpha<$ 1 and $C>0$ depending only on $n,\lambda,\Lambda$, $\operatorname{diam}G$, $C_{0}$ and $C_{1}$, such that for any $Q(X_{1},R)\subset G$, $r\leq R$,
$\operatorname{osc}_{Q(X_{1},r)}Dv\leq C\Big{(}\frac{r}{R}\Big{)}^{\alpha}(\operatorname{osc}_{Q(X_{1},R)}Dv+R),$
where $Q(X_{1},R)=\{(x,t):|x-x_{1}|<R,t_{1}-R^{2}<t\leq t_{1}\}$ with $X_{1}=(x_{1},t_{1})$.
This lemma is derived from Theorem 12.3 in [20]. Note that this lemma reveals the interior $\tilde{C}^{1,\alpha}$ estimate of $v$. The next lemma is the weak Harnack inequality for linear parabolic equations, borrowed from Lemma 7.37 in [20].
Lemma 2.10.
Under the same conditions as in Lemma 2.9. Suppose also that $v$ is nonnegative. Then for any $Q(X_{1},4R)\subset G$, we have
[…]
Here $\Theta(X_{1},R)=Q((x_{1},t_{1}-4R^{2}),R)$ and $p,C$ are positive constants depending only on $n,\lambda,\Lambda$.
Additionally, the next lemma is Lemma 12.4 in [20].
Lemma 2.11.
Let $w\in C(\overline{Q})$ and $Dw\in C(Q)$ for some cylinder $Q=Q(X_{0},R)$. Suppose that there are nonnegative constants $\bar{C}_{1},\bar{C}_{2}$ and $\alpha<1$ such that
$\operatorname{osc}\limits_{Q(X_{1},r)}Dw\leq\bar{C}_{1}\Big{(}\frac{r}{\mu}\Big{)}^{\alpha}(\operatorname{osc}\limits_{Q(X_{1},\mu)}Dw+\bar{C}_{2}\mu^{\alpha}),$
whenever $Q(X_{1},\mu)\subset Q$ and $r\leq\mu$. Then for any $\ell\in\mathbb{R}^{n}$ and $a\in\mathbb{R}$, we have
$\sup_{Q(X_{0},R/2)}|Dw-\ell|\leq C(\bar{C}_{1},n,\alpha)\Big{(}R^{-1}\sup_{Q}|w(X)-\ell\cdot x-a|+\bar{C}_{2}R^{\alpha}\Big{)}.$
We will use Lemma 2.10 to prove an oscillation estimate for $\frac{v}{x_{n}}$, which gives the Hölder estimate for $v_{n}$ on $F$, the flat boundary. Combining this with the interior estimates, Lemma 2.9, and with the help of Lemma 2.11, we will derive a global $\tilde{C}^{1,\alpha}$ estimate for $v$ in $\frac{1}{2}G:=B_{R_{0}/2}^{+}\times(t_{0}-R_{0}^{2}/4,t_{0}]$. Recall the definition of $v$, then the global Hölder estimate of $D^{2}u$ on $\Omega\times(\epsilon,T]$ follows from equation (2.1).
Theorem 2.12.
Let $v\in\tilde{C}^{2}(\overline{G})$ be a solution of equation (2.22) with conditions (2.23) and (2.24). Then there are small positive constants $r_{0}$ and $\delta$ depending on $n,\lambda,\Lambda,R_{0}$ and $C_{0}$ such that for all $r\in(0,r_{0})$ and $X_{1}\in Q_{r,\delta}=\{(x^{\prime},x_{n},t):|x^{\prime}|<r,0<x_{n}<\delta r,t_{0}-r^{2}<t\leq t_{0}\},$ we have
$|Dv(X_{1})-Dv(0,t_{0})|\leq Cr^{\alpha},$
for some positive constants $C$ and $\alpha<1$ depending on $n,C_{0},C_{1},\lambda,\Lambda,r_{0},\delta$ and $R_{0}$.
Proof.
Take $r<r_{0}$ small such that $r_{0}+r_{0}/\delta\leq 1/2R_{0}$. $\delta<1$ is a small constant to be chosen later. For any $X_{1}=(x_{1},t_{1})\in Q_{r,\delta}$, let $\ell=Dv(X_{1}^{\prime})$, where $X_{1}^{\prime}=(x_{1}^{\prime},0,t_{1})=(x_{1,1},\cdots,x_{1,n-1},0,t_{1})\in F$. Since $v=0$ on $F$, $\ell=(0,\cdots,0,\ell_{n})$. Take any point
$X=(x,t)\in Q(X_{1},\mu)=\{|x-x_{1}|<\mu,t_{1}-\mu^{2}<t\leq t_{1}\}\quad\text{with $\mu=\frac{1}{2}x_{1,n}$}.$
If we take $\sigma=r+\mu/\delta$, then $\sigma<2r$ and $X,X_{1}^{\prime}\in Q_{\sigma,\delta}\subset Q_{R_{0}/2,\delta}$. Consider the function $v(X)-\ell\cdot x$, then by Lemma 2.9 and Lemma 2.11,
(2.25) $\begin{split}|Dv(X_{1})-Dv(X_{1}^{\prime})|&\leq\sup_{Q(X_{1},\mu/2)}|Dv-\ell|\\ &\leq C\Big{(}\mu^{-1}\sup_{Q(X_{1},\mu)}|v(X)-\ell\cdot x|+\mu^{\alpha}\Big{)}.\end{split}$
We claim that
(2.26) $\operatorname{osc}\limits_{Q_{\sigma,\delta}}\frac{v(X)}{x_{n}}\leq C\sigma^{\alpha}.$
Then since $v|_{F}=0$ and $X,X_{1}^{\prime}\in Q_{\sigma,\delta}$, we get
$\Big{|}\frac{v(X)}{x_{n}}-\ell_{n}\Big{|}\leq\operatorname{osc}\limits_{Q_{\sigma,\delta}}\frac{v(X)}{x_{n}}\leq C\sigma^{\alpha}.$
Thus (2.25) gives
$|Dv(X_{1})-Dv(X_{1}^{\prime})|$ $\leq C\Big{(}\mu^{-1}\sup_{Q(X_{1},\mu)}Cx_{n}\sigma^{\alpha}+\mu^{\alpha}\Big{)}$
$\leq C(\mu^{-1}\sigma^{\alpha}(x_{1,n}+\mu)+\mu^{\alpha})$
$\leq C(\sigma^{\alpha}+\mu^{\alpha})$
$\leq Cr^{\alpha},$
where we use the fact that $\mu=\frac{1}{2}x_{1,n}<\delta r$ and $\sigma<2r$. Again by the claim,
$|Dv(X_{1}^{\prime})-Dv(0,t_{0})|\leq C\sigma^{\alpha}\leq Cr^{\alpha},$
then we have proved the theorem by the triangle inequality. For the rest part of this proof, we will show the claim (2.26).
First suppose that $v\geq 0$ and denote
$F_{R,\delta}=:\{(x,t):|x^{\prime}|<R,x_{n}=\delta R,t_{0}-R^{2}<t\leq t_{0}\}.$
We will prove the following inequality
(2.27) $\inf_{F_{R,\delta}}\frac{v}{x_{n}}\leq 10\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}\quad\text{ for $R<1/4R_{0}$.}$
Without loss of generality, assume that
(2.28) $R=1,\;\inf_{F_{1,\delta}}\frac{v}{x_{n}}=1.$
Let
$w(x,t)=\left[1-|x^{\prime}|^{2}+(1+\sup|f|)\frac{x_{n}-\delta}{\sqrt{\delta}}+t-t_{0}\right]x_{n},$
then for $L=-\partial_{t}+a^{ij}\partial_{ij}$, we have
$Lw\geq-\delta+\frac{2(1+\sup|f|)}{\sqrt{\delta}}\lambda-c(n)\Lambda(1+\delta)\quad\text{in $Q_{1,\delta}$.}$
Hence, $Lw\geq\sup|f|\geq Lv$ in $Q_{1,\delta}$, when $\delta=\delta(\lambda,\Lambda,c(n))>0$ is small enough. On $\{x_{n}=0\}\cap\partial_{p}Q_{1,\delta}$, $w=0=v$. On $\{x_{n}=\delta\}\cap\partial_{p}Q_{1,\delta}$, $w\leq\delta(1-|x^{\prime}|^{2})\leq\delta\leq v$ by (2.28). Additionally, on $\{|x^{\prime}|=1\}\cap\partial_{p}Q_{1,\delta}$ and $\{t=t_{0}-1\}\cap\partial_{p}Q_{1,\delta}$, $w<0\leq v$. Thus by maximal principle, we have $w\leq v$ in $Q_{1,\delta}$. Therefore, in $Q_{1/2,\delta}$,
$\frac{v}{x_{n}}$ $\geq 1-|x^{\prime}|^{2}+(1+\sup|f|)\frac{x_{n}\delta}{\sqrt{\delta}}+t-t_{0}$
$\geq 1-\frac{1}{4}-\sqrt{\delta}(1+\sup|f|)-\frac{1}{4}$
$\geq\frac{1}{10}-\sup|f|.$
when $\delta<1/100$. Now the constant $\delta$ is determined, so does $r_{0}$.
Next, without the assumption (2.28), let
$m_{1}=\inf_{F_{R,\delta}}\frac{v}{x_{n}},\quad\text{and}\quad\tilde{v}(x,t)=\frac{v(Rx,R^{2}(t-t_{0})+t_{0})}{m_{1}R}.$
Then $-\tilde{v}_{t}+a^{ij}\tilde{v}_{ij}=\tilde{f}$ in $Q_{1,\delta}$, with $\tilde{f}=\frac{R}{m_{1}}f$. Note that $\inf_{F_{1,\delta}}\frac{\tilde{v}}{x_{n}}=1$. Then from the above arguments,
$\frac{\tilde{v}(x,t)}{x_{n}}\geq 1/10-\sup|\tilde{f}|.$
Hence in the original coordinates, we have
$\frac{v}{x_{n}}\geq\frac{m_{1}}{10}-R\sup|f|\quad\text{in $Q_{R/2,\delta}$,}$
which yields (2.27).
Denote
$Q_{R,\delta}^{*}=\{|x^{\prime}|<R,\frac{1}{2}\delta R<x_{n}<\delta R,t_{0}-R^{2}<t\leq t_{0}\},$
$Q^{\prime}_{R,\delta}=\{|x^{\prime}|<R,\frac{1}{2}\delta R<x_{n}<\delta R,t_{0}-3R^{2}<t\leq t_{0}-2R^{2}\},$
and
$m=\inf_{Q_{2R,\delta}}\frac{v}{x_{n}},\quad\quad M=\sup_{Q_{2R,\delta}}\frac{v}{x_{n}}.$
Then $Mx_{n}-v\geq 0$, $v-mx_{n}\geq 0$ in $Q_{2R,\delta}$. Applying Lemma 2.10, the weak Harnack inequality, to $Mx_{n}-v$ and $v-mx_{n}$ in $Q^{*}_{R,\delta}$, we have
(2.29) $\begin{split}\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(Mx_{n}-v)^{p}\Big{)}^{1/p}&\leq C\Big{(}\inf_{Q^{*}_{R,\delta}}(Mx_{n}-v)+R^{2}\sup|f|\Big{)}\\ &\leq C(R\inf_{F_{R,\delta}}\Big{(}M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq C\Big{(}R\inf_{Q_{R/2,\delta}}(M-\frac{v}{x_{n}})+R^{2}\sup|f|\Big{)}\\ &\leq CR\Big{(}M-\sup_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}.\end{split}$
The third inequality above is deduced by (2.27). Similarly we have
(2.30) $\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}(v-mx_{n})^{p}\Big{)}^{1/p}\leq CR\Big{(}\inf_{Q_{R/2,\delta}}\frac{v}{x_{n}}-m+R\sup|f|\Big{)}.$
Add equations (2.29) and (2.30) up, and by Minkowski inequality,
$(M-m)\Big{(}R^{-n-2}\int_{Q^{\prime}_{R,\delta}}x_{n}^{p}\Big{)}^{1/p}\leq CR\Big{(}M-m-\operatorname{osc}\limits_{Q_{R/2,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}.$
Direct computations give that
$\operatorname{osc}\limits_{Q_{R/2,\delta}}\frac{v}{x_{n}}\leq C\Big{(}\operatorname{osc}\limits_{Q_{2R,\delta}}\frac{v}{x_{n}}+R\sup|f|\Big{)}.$
Then it follows from the classical iterative arguments (see Lemma 8.23 in [10] for a reference) that
(2.31) $\operatorname{osc}\limits_{Q_{R,\delta}}\frac{v}{x_{n}}\leq C\Big{(}\frac{R}{R_{0}}\Big{)}^{\beta}\Big{(}\operatorname{osc}\limits_{Q_{R_{0},\delta}}\frac{v}{x_{n}}+R_{0}\sup|f|\Big{)}$
for any $R<R_{0}$ and some $\beta<1$ depending on $n,C_{0},C_{1},\lambda,\Lambda,\delta$. We still use $\alpha$ to denote $\min\{\alpha,\beta\}$. Since $|v|+|Dv|$ and $|f|$ are bounded in $\overline{G}$, the claim (2.26) is now proved. ∎
Since $v|_{F}=0$, the inequality (2.31) implies the Hölder continuity of $Dv$ on $F$. With the above theorem, we infer that,
$|Dv(X)-Dv(Y)|\leq C(|x-y|+|t-s|^{1/2})^{\alpha}$
for any $X\in G$ and $Y\in F$. The $\tilde{C}^{\alpha}$ estimate for $Dv$ on $\overline{G}$ then follows from the triangle inequality and Lemma 2.9.
Reverting to $u$, we have
Corollary 2.13.
For any $\epsilon>0$, there exist positive constants $C$ and $\alpha$ such that
$||D^{2}u||_{\tilde{C}^{\alpha}(\overline{\Omega}\times(\epsilon,T])}\leq C,$
where the constants $C$ and $\alpha$ depend only on $\epsilon^{-1}$ and those universal quantities in Theorem 2.3.
Proof.
Recall that $v(x,t)=T_{k}(u-\phi)(x^{\prime},x_{n}+\rho(x^{\prime}),t)$ in $\Psi(B_{\rho_{0}}(Y_{0}))$ with $T_{k}=\partial_{y_{k}}+\rho_{y_{k}}\partial_{y_{n}}$, $1\leq k<n$, where we take $\rho_{0}>0$ small enough. If we define $\tilde{u}(x,t)=u(x^{\prime},x_{n}+\rho(x^{\prime}),t)$, then $\tilde{u}_{ij}=T_{i}T_{j}u$, for any $1\leq i,j\leq n$. What we have proved above is equivalent to
(2.32) $||\tilde{u}_{kl}||_{\tilde{C}^{\alpha}(\overline{G})}\leq C\text{ for any }1\leq k<n\text{ and }1\leq l\leq n.$
To prove the $\tilde{C}^{\alpha}$ estimate for $\tilde{u}_{nn}$, note that the equation (2.1) can be written as $\det(P_{i}P_{j}\tilde{u})=\tilde{u}_{t}+\psi(\Psi^{-1}(x),t)$, with $P_{k}=\partial_{x_{k}}-\rho_{x_{k}}\partial_{x_{n}}$ for $1\leq k<n$ and $P_{n}=\partial_{x_{n}}$, which is equivalently
$Q_{1}\tilde{u}_{nn}+Q_{2}=\tilde{u}_{t}+\psi(\Psi^{-1}(x),t),$
where $Q_{1}$ and $Q_{2}$ are polynomials of $\tilde{u}_{kl}$ with $k+l<2n$ and $\rho$ up to the second derivatives. By Theorem 2.5, the eigenvalues of the Hessian $D^{2}u$ have positive lower and upper bounds; therefore, $Q_{1}$ has a positive lower bound. This fact together with Theorem 2.7 and (2.32) leads to the $\tilde{C}^{\alpha}$ estimate of $\tilde{u}_{nn}$. From this the $\tilde{C}^{\alpha}$ estimate of $D^{2}u$ in $\Omega\times(\epsilon,T]$ follows by the arbitrariness of $Y_{0}$ and the interior higher regularity of $u$ in Theorem 2.6. ∎
3. Existence and uniqueness
Equation (2.1) is uniformly parabolic by virtue of Theorem 2.5. Hence with some compatibility conditions on $\phi$, we have the following short-time existence theorem by the inverse function theorem. The proof is a modification of the argument in [9, Theorem 2.5.7].
Theorem 3.1.
Let $\Omega$ be a uniformly convex domain in $\mathbb{R}^{n}$ with smooth boundary. Assume that $\varphi$ is a smooth and uniformly convex function on $\overline{\Omega}$, $\psi$ is a smooth function on $\overline{\Omega}\times[0,T]$ and $h$ is a smooth function on $\partial\Omega\times[0,T]$. Suppose also that the equation
[…]
satisfies the compatibility condition of order $1$, i.e.,
$h(0)=\varphi|_{\partial\Omega}\quad\text{ and }\quad-h_{t}(0)+\det D^{2}\varphi|_{\partial\Omega}=\psi(\cdot,0)|_{\partial\Omega}.$
Then for any $0<\alpha\leq 1$, there is a small positive constant $\epsilon$ such that the initial boundary value problem (3.1) has a unique solution $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$, which is uniformly convex in $x$ for all $t\in[0,2\epsilon)$. Here we have $||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})}\leq C$ and $Q_{2\epsilon}=\Omega\times(0,2\epsilon]$. The constants $\epsilon$ and $C$ depend on $n$, $\Omega$, the uniform convexity of $\varphi$, $||\psi||_{\tilde{C}^{1}(\overline{\Omega}\times[0,T])}$, $||\varphi||_{\tilde{C}^{3}(\overline{\Omega})}$, and $||h||_{\tilde{C}^{2}(\partial\Omega\times[0,T])}$.
Proof.
Consider the equation
[…]
According to theorem 5.2 in Chapter 4 of [19], problem (3.2) has a unique solution $w\in\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$ for some positive constant $T_{0}$. Moreover, $w(x,t)$ is still uniformly convex in $x$ for any $0\leq t\leq T_{0}$ when $T_{0}$ is small enough. Indeed, by the $\tilde{C}^{2,\alpha}$ regularity of $w$, $||w(\cdot,t)-\varphi||_{C^{2}(\overline{\Omega})}\leq Ct^{\alpha/2}$, then the uniform convexity of $w(\cdot,t)$ follows from that of $\varphi$ for sufficiently small $t$.
Let
$g=-w_{t}+\det D^{2}w-\psi\in\tilde{C}^{\alpha}(\overline{Q}_{T_{0}}).$
Then by equation (3.2) and the compatibility condition, $g|_{t=0}=0$. Let $V$ be an open neighborhood of $w$ in $\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$. Define the map $\Phi:V\to W$ by
(3.3) $\Phi(u)=(-u_{t}+\det D^{2}u-\psi,~{}~{}u|_{t=0},~{}~{}u|_{\partial\Omega\times[0,T_{0}]}),$
where $W=\tilde{C}^{\alpha}(\overline{Q}_{T_{0}})\times S$ and
$S=\{(a,b)\in C^{2,\alpha}(\overline{\Omega})\times\tilde{C}^{2,\alpha}(\partial\Omega\times[0,T_{0}]):\;a|_{\partial\Omega}=b|_{t=0},\det D^{2}a|_{\partial\Omega}=(\psi+b_{t})|_{t=0}\}.$
One can prove that $V$ and $W$ are Banach manifolds (see theorem 4.4 in [15]). Note that $\Phi(w)=(g,\varphi,h)$. The differential of $\Phi$ at $w$ is a linear operator denoted by $L$.
$L\eta=(-\eta_{t}+(\det D^{2}w)w^{ij}\eta_{ij}-\psi_{t},~{}~{}\eta(0),~{}~{}\eta|_{\partial\Omega\times[0,T_{0}]}),\quad\text{for any $\eta\in V$.}$
The construction of $w$ and $S$ ensures the compatibility condition for $L$. Hence by Theorem 5.2 in Chapter 4 of [19], $L$ is an isomorphism from the tangent space of $V$ at $w$ to that of $W$ at $(g,\varphi,h)$. Then by the inverse function theorem [5, Section 15], $\Phi$ is an $C^{1}$ diffeomorphism from $B_{\tilde{\rho}}(w)$, a small neighborhood of $w$ in $\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$ to an open neighborhood $U$ of $(g,\varphi,h)$ in $W$. Take $\tilde{\rho}$ small enough, then the functions in $B_{\tilde{\rho}}(w)$ is still uniformly convex in $x$.
Let $\epsilon\in(0,T_{0})$ be small, $\eta_{\epsilon}\in C^{\infty}([0,1])$ such that $0\leq\eta_{\epsilon}\leq 1$, $0\leq\eta_{\epsilon}^{\prime}\leq C\epsilon^{-1}$, and $\eta_{\epsilon}=0$, if $0\leq t\leq 2\epsilon$; $\eta_{\epsilon}=1$, if $4\epsilon\leq t\leq 1$. Denote $g_{\epsilon}=g\eta_{\epsilon}$. Then $g_{\epsilon}\in\tilde{C}^{\alpha}(\overline{Q}_{T_{0}})$. One can prove that $g_{\epsilon}\to g\text{ in }\tilde{C}^{\alpha}(\overline{Q}_{T_{0}})$ when $\epsilon\to 0$. Thus by taking $\epsilon$ small enough, $(g_{\epsilon},\varphi,h)\in U\subset W$, and there is a function $u\in B_{\tilde{\rho}}(w)\subset\tilde{C}^{2,\alpha}(\overline{Q}_{T_{0}})$ such that
$\Phi(u)=(g_{\epsilon},\varphi,h).$
Since $g_{\epsilon}=0$ for $0\leq t\leq 2\epsilon$, then $u\in\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})$ is a solution to equation (3.1) in $\Omega\times(0,2\epsilon]$ and is uniformly convex in $x$. ∎
Remark 3.2.
When the initial-boundary condition of equation (3.1) satisfies the compatibility condition of order $m=[\frac{l+\alpha}{2}]+1$, there exists a unique solution $u\in\tilde{C}^{l+2,\alpha}(\overline{Q}_{\epsilon})$. The proof is almost the same as above, and here we omit it. One can refer to Chapter 4, Section 5 in [19] for the general definition of $m$-th order compatibility condition. To be more explicit, the compatibility condition of m-th order for our equation means that
$u^{(k)}(x)|_{x\in\partial\Omega}=\frac{\partial^{k}h}{\partial t^{k}}\Big{|}_{t=0},$
for $k=0,\cdots,m$, where $u^{(k)}$ is recursively defined by
$u^{(0)}(x)=\varphi(x),\quad u^{(k+1)}(x)=\frac{\partial^{k}}{\partial t^{k}}\Big{|}_{t=0}\Big{(}\det D^{2}u(x,t)-\psi(x,t)\Big{)}.$
Here the right term of the second equality can be recursively expressed by derivatives of $\varphi$ and $\psi$.
The uniqueness of solutions to the equation (2.1) then follows immediately from Lemma 2.1.
Corollary 3.3.
Let $\Omega$ be a uniformly convex bounded domain with smooth boundary and $\phi$ be a continuous function in $\overline{Q}_{T}$. Then equation (2.1) admits at most one solution in $\tilde{C}^{2}$ that is convex in $x$.
3.1. Proof of main theorems
We are in a position to prove our main theorems.
Proof of Theorem 1.2:  By uniqueness, the a priori solution $u$ in Corollary 2.13 coincides with the solution derived from Theorem 3.1 on $[0,2\epsilon]$ if $\varphi=\phi|_{t=0}$ and $h=\phi|_{\partial\Omega}$. Thus, for some small positive constant $\epsilon$, $\|u\|_{\tilde{C}^{2,\alpha}(\overline{Q}_{2\epsilon})}$ is bounded. Combined with Corollary 2.13, this gives the global $\tilde{C}^{2,\alpha}$ estimate to $u$ on $\overline{Q}_{T}$. ∎
Proof of Theorem 1.3:  The existence of $\tilde{C}^{2,\alpha}$ solutions of equation (2.1) can be derived step by step from the short-time existence under the uniform global $\tilde{C}^{2,\alpha}$ estimate. It is also unique by Corollary 3.3.
Additionally, the higher order regularity of $u$ can also be obtained by classic regularity theory of linear parabolic equations iteratively if $\partial\Omega$ and $\phi$ are smooth enough, and equation (2.1) admits such solution uniquely if the equation satisfies higher order compatibility conditions in Remark 3.2. Specifically, let $k\geq 2$, $0<\beta\leq 1$. Assume that $\partial\Omega$ is of class $C^{k+2,\beta}$, $\phi\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$, and equation (2.1) satisfies the compatibility condition of order $[\frac{k+\beta}{2}]+1$. Then the equation admits a unique solution $u\in\tilde{C}^{k+2,\beta}(\overline{Q}_{T})$ that is convex in $x$. ∎
3.2. Necessity of condition $(\mathcal{P}3)$
Conditions $({\mathcal{P}1})$, $({\mathcal{P}2})$ and $({\mathcal{P}3})$ were introduced by Ivochkina and Ladyzhenskaia [14] to study (1.1). Note that conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ are natural requirements to ensure $C^{-1}\leq\det D^{2}u\leq C$ and thus guarantee the convexity of $u$. We will take counterexamples to show that condition $({\mathcal{P}3})$ is also necessary for the existence of convex solutions to (2.1).
In one space dimension, take $\Omega=(0,1)$ and $T>1$. Then equation (2.1) reduces to the following initial-boundary value problem
{$&-v_{t}+v_{xx}=\psi(x,t)&&\text{ in }Q_{T}=(0,1)\times(0,T],\\ &v=\phi&&\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $\psi$ and $\phi$ are smooth functions satisfying conditions $({\mathcal{P}1})-({\mathcal{P}3})$. Assume $v$ is a smooth solution to this equation that is convex in $x$. Define a smooth bump function $w:\overline{Q}_{T}\to\mathbb{R}$ by
(3.4) $w(x,t)=\left\{\begin{aligned} &At\cdot e^{\frac{-B}{x^{2}(1-x)^{2}}}&&\text{ in }Q_{T},\\ &0&&\text{ on }\partial_{p}Q_{T},\end{aligned}\right.$
where $A,B$ are positive constants. The function $w$ is non-negative, smooth up to the boundary, and satisfies that
(3.5) $w_{t}=0\ \text{on}\ \partial\Omega\times[0,T];\quad\frac{\partial^{k}w}{\partial x^{k}}=0\ \text{on}\ \partial_{p}Q_{T}\ \ \text{for}\ k=0,1,2,\cdots.$
Set $u=v+w$; then we have
(3.6) {$&-u_{t}+u_{xx}=\psi+(-w_{t}+w_{xx})=:\psi+\rho&&\text{ in }Q_{T},\\ &u=\phi&&\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
Since (3.5) gives that $\rho=0$ on $\partial\Omega\times[0,T]$, then conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ are satisfied for $\phi$ and $\psi+\rho$. Direct computation yields
$w_{t}=Ae^{\frac{-B}{x^{2}(1-x)^{2}}},\quad w_{xx}=-2ABt\cdot e^{\frac{-B}{x^{2}(1-x)^{2}}}\frac{P_{6}(x)}{x^{6}(x-1)^{6}}.$
Here $P_{6}(x)$ is a sixth-degree polynomial given by
$P_{6}(x)=2B-8Bx+(8B-3)x^{2}+16x^{3}-33x^{4}+30x^{5}-10x^{6}.$
One can check that there exist constants $x_{0}\in(0,1/2)$ and $c_{1},c_{2}>0$ depending only on $B$ such that
$\inf_{x\in(0,1)}\rho(x,1)=\rho(x_{0},1)=-A\cdot c_{1}<0,\quad\sup_{x\in(0,1)}\rho(x,1)=\rho(1/2,1)=A\cdot c_{2}>0.$
Then by choosing $A$ sufficiently large, $\rho$ attains both a negative minimum and a positive maximum, each with large absolute values. This destroys the concavity of $\psi+\rho$. For any fixed $t$, the maximum of $w(\cdot,t)$ on $(0,1)$ is $At\cdot e^{-16B}$. Then, the maximum of $w$ can be made arbitrarily large and $u$ fails to remain convex in $x$; therefore, by the comparison principle, equation (3.6) does not admit any solution that is convex in $x$ for all time.
In $n$ space dimension, a counterexample can also be constructed similarly for rotationally symmetric solutions. Let $v(r,t)$ be a smooth radial solution on $(0,1)\times(0,T]$ satisfying
{$&-v_{t}+(v_{r}/r)^{n-1}v_{rr}=\psi(r,t)&&\text{ in }Q_{T}=(0,1)\times(0,T],\\ &v=\phi&&\text{ on }\partial_{p}Q_{T}.\end{aligned}\right.$
And $\bar{v}(x,t)=v(|x|,t)$ solves the parabolic Monge–Ampère equation
{$&-\bar{v}_{t}+\det D^{2}\bar{v}=\psi(|x|,t)&&\text{ in }U_{T}=B_{1}(0)\times(0,T],\\ &\bar{v}=\phi(|x|,t)&&\text{ on }\partial_{p}U_{T}.\end{aligned}\right.$
We use the same one‐dimensional bump $w(r,t)$ as in (3.4) above and set $\bar{w}(x,t)=w(|x|,t)$. The equation for $\bar{u}=\bar{v}+\bar{w}$ becomes
(3.7) {$&-\bar{u}_{t}+\det D^{2}\bar{u}=\bar{\Psi}(x,t)&&\text{in }U_{T},\\ &\bar{u}=\phi(|x|,t)&&\text{ on }\partial_{p}U_{T},\end{aligned}\right.$
where
$\bar{\Psi}(x,t)=\Psi(|x|,t)=\Psi(r,t)=-v_{t}-w_{t}+\big{(}\frac{v_{r}+w_{r}}{r}\big{)}^{n-1}v_{rr}+\big{(}\frac{v_{r}+w_{r}}{r}\big{)}^{n-1}w_{rr}.$
By (3.5), $\bar{\Psi}$ coincides with $\psi$ on $\partial B_{1}(0)\times[0,T]$, and $\bar{u}$ matches $\bar{v}$ on $\partial_{p}U_{T}$. Thus, conditions $({\mathcal{P}1})$ and $({\mathcal{P}2})$ hold for arbitrary $A$ and $B$. For sufficiently large $A$, there exist $r_{0}\in(1/2,1)$ and $c_{3},c_{4}>0$ depending on $B$ such that $w_{r}(r_{0},1)=A\cdot c_{3}>0$ and $w_{rr}(r_{0},1)=-A\cdot c_{4}<0$. Since $||v||_{\tilde{C}^{2}(\overline{Q}_{T})}$ is bounded, $\Psi(r_{0},1)$ can be made arbitrarily negative, which destroys the concavity of $\Psi$ in $r$. Additionally, $u$ loses convexity in $r$ when $w$ attains a sufficiently large maximum. From the comparison principle Lemma 2.1, we know that there is no convex solution to equation (3.7).
4. Gauss curvature flow
In this section, we consider the Gauss curvature flow with boundary
(4.1) {$u_{t}&=\frac{\det D^{2}u}{(1+|Du|^{2})^{\frac{n+1}{2}}}\quad\ \text{ in }Q_{T},\\ u&=\phi\quad\quad\quad\quad\quad\quad\quad\,\;\,\text{on }\partial_{p}Q_{T}.\end{aligned}\right.$
where $\phi$ and $Q_{T}$ are defined as before. We will obtain the a priori estimate for equation (4.1) using the same method as before.
Theorem 4.1.
Let $u\in\tilde{C}^{4}(\overline{Q}_{T})$ be convex in $x$ and solve equation (4.1). Suppose that $\Omega$ is a uniformly convex domain in $\mathbb{R}^{n}$, $\partial\Omega\in C^{4}$, $\phi\in\tilde{C}^{4}(\overline{Q}_{T})$, $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$ and $\phi(\cdot,0)$ is uniformly convex. Then we have
$||u||_{\tilde{C}^{2,\alpha}(\overline{Q}_{T})}\leq C$
for some positive constants $C$ and $\alpha$ depending only on $n,T,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
Proof.
The proof of this theorem is similar to that of Theorem 1.2, except that the linearized equation of (4.1) has lower order terms concerning $Du$.
(a) $C^{0}$ and $C^{1}$ estimates. Note that $u$ reaches its maximum and minimum value on $\partial_{p}Q_{T}$, then $||u||_{C^{0}(\overline{Q}_{T})}\leq||\phi||_{C^{0}(\overline{Q}_{T})}\leq C$.
For convenience we denote $\beta=\frac{n+1}{2}$. Differentiating the equation (4.1), we have
$\mathcal{L}^{\prime}u_{k}=0,\quad\text{for}\ \ k=1,\cdots,n,\quad\text{and}\quad\mathcal{L}^{\prime}u_{t}=0$
with
(4.2) $\mathcal{L}^{\prime}$ $=-\partial_{t}+\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta}}u^{ij}\partial_{ij}-2\beta\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta+1}}u_{i}\partial_{i}$
$=-\partial_{t}+u_{t}u^{ij}\partial_{ij}-\frac{2\beta}{1+|Du|^{2}}u_{t}u_{i}\partial_{i}~{}.$
Then $\mathcal{L}^{\prime}$ is a linear parabolic operator. It follows that $|u_{t}|$ is bounded by $\sup_{\partial_{p}Q_{T}}|u_{t}|$ and
$C^{-1}\leq u_{t}\leq C$
for some positive constant $C$ depending on $n,c_{0},\Omega$, $||\phi||_{\tilde{C}^{2}}$ and the uniform convexity of $\phi(\cdot,0)$.
To prove the bound of $|Du|$, we apply the similar auxiliary function in [24]. Fix any $t_{0}\in[0,T]$. Let $\tilde{\phi}$ be the solution of the equation
$\det D^{2}\tilde{\phi}&=1\quad\quad\quad\quad\text{in }\Omega,\\ \tilde{\phi}&=\phi(\cdot,t_{0})\quad\,\text{ on }\partial\Omega.\end{aligned}\right.$
The existence and $C^{2}$ bound of solutions to this equation are guaranteed by [1]. We consider the barrier function
(4.3) $\overline{w}=\tilde{\phi}-\frac{1}{q}\log(1+kd_{\partial\Omega})$
defined in $\Omega_{\delta}=\{x\in\Omega:\,d_{\partial\Omega}(x)<\delta\}$ with small positive constant $\delta$ such that $\overline{w}\in C^{4}(\Omega_{\delta})$.
The positive constants $\delta,q$, and $k$ shall be determined later. Direct computations give that
$\frac{\det D^{2}\overline{w}}{(1+|D\overline{w}|^{2})^{\beta}}$ $\geq\frac{1+Ck^{n+1}q^{-n}(1+kd_{\partial\Omega})^{-n-1}}{(1+2|D\tilde{\phi}|^{2}+2k^{2}q^{-2}(1+kd_{\partial\Omega})^{-2})^{\beta}}$
$\geq Cq,$
when $k^{2}q^{-2}(1+k\delta)^{-2}\geq\max_{\overline{\Omega}}(1+2|D\tilde{\phi}|^{2})$ and $\delta$ is small enough.
We first fix the sufficiently big constant $q$ such that $Cq\geq\max u_{t}$. Thus
$\frac{\det D^{2}\overline{w}}{(1+|D\overline{w}|^{2})^{\beta}}\geq\frac{\det D^{2}u}{(1+|Du|^{2})^{\beta}}\quad\text{ in }\Omega_{\delta}\times\{t_{0}\}.$
Then we take $\delta$ to be small enough and $k\delta\geq 1$ to make
$k^{2}q^{-2}(1+k\delta)^{-2}\geq\frac{1}{4}q^{-2}\delta^{-2}\geq\max_{\overline{\Omega}}(1+2|D\tilde{\phi}|^{2}).$
Finally, we take $k$ sufficiently large to satisfy
$\log(k\delta)\geq q(\max_{\overline{\Omega}}|\tilde{\phi}|+\max_{\overline{Q}_{T}}|u|),$
then $\overline{w}\leq u$ on $\partial\Omega_{\delta}\times\{t_{0}\}$. Therefore $u\geq\overline{w}$ in $\Omega_{\delta}\times\{t_{0}\}$ by the comparison principle. Then for the exterior normal vector $\nu$ at any point $x_{0}$ on $\partial\Omega$, we have $u_{\nu}\leq w_{\nu}\leq C$. By convexity of $u$ in $x$ and arbitrariness of $t_{0}$, this proves the bound for $|Du|$ on $\overline{Q}_{T}$.
(b) $C^{2}$ estimates on the boundary. We can prove this by the same method as before but with different barrier functions. As for the second tangential derivatives, we can also regard $u_{t}(1+|Du|^{2})^{\frac{n+1}{2}}$ which has positive lower and upper bounds as a data function. Then it follows that
(4.4) $C^{-1}I_{n-1}\leq D_{x^{\prime}x^{\prime}}u\leq CI_{n-1}\quad\text{on}\quad\partial_{p}Q_{T}.$
To prove the uniform bound for the mixed second derivatives, we use the comparison principle for $\mathcal{L}^{\prime}$. Denote $\mathcal{T}$ as in (2.13). Then
(4.5) $|\mathcal{L}^{\prime}\mathcal{T}(u-\phi)|\leq\underline{a}\Big{(}1+\sum_{i=1}^{n}u^{ii}\Big{)}$
for some positive constant $\underline{a}$ depending on $||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$, the positive lower and upper bounds for $u_{t}$ and the bound for $|Du|$.
Recall that $0\in\partial\Omega$ and the boundary $\partial\Omega$ can be represented as in (2.12) near 00. On $\Omega\times\{0\}$, $\mathcal{T}(u-\phi)=0$ and
(4.6) $|\mathcal{T}(u-\phi)|\leq\bar{a}|x^{\prime}|^{2},\quad\text{on $(B_{\varepsilon}\cap\partial\Omega)\times[0,T]$}.$
In $(B_{\varepsilon}\cap\Omega)\times[0,T]$ with $B_{\varepsilon}=\{|x|<\varepsilon\}$, we shall consider the following barrier function
$\zeta(x,t)=\frac{1}{2}\sum_{\alpha,\beta<n}(B_{\alpha\beta}-\mu\delta_{\alpha\beta})x_{\alpha}x_{\beta}+\frac{1}{2}Mx_{n}^{2}-Nx_{n},$
where $\mu$ is a fixed small constant such that $B_{\alpha\beta}-\mu\delta_{\alpha\beta}>0$, and $M,N$ will be determined later.
First we take $\varepsilon$ small, $M\varepsilon\leq 2$ and $N$ larger than $1+\lambda_{max}/c$, where $\lambda_{max}$ is the maximum eigenvalue of the $(n-1)\times(n-1)$ matrix $(B_{\alpha\beta}-\mu\delta_{\alpha\beta})$, and $c$ is the constant such that $x_{n}\geq c|x^{\prime}|^{2}$ on $\partial\Omega$. Then we have $\zeta<0$ in $(B_{\varepsilon}\cap\Omega)\times\{0\}$.
Thus
(4.7) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in $(B_{\varepsilon}\cap\Omega)\times\{0\}$ for all positive constant $A$.}$
Note that we can choose a large $M\geq 2$ and a small $c_{1}>0$ such that
(4.8) $\begin{split}\mathcal{L}^{\prime}\zeta&\geq c_{1}\Big{(}\sum_{i=1}^{n}u^{ii}+Mu^{nn}\Big{)}-c_{2}\\ &\geq\frac{1}{2}c_{1}\sum_{i=1}^{n}u^{ii}-c_{2}+c_{3}M^{1/n}.\end{split}$
Here we use the bounds for $u_{t}$ and the fact that
$\begin{split}\frac{1}{n}\Big{(}\sum_{i=1}^{n}u^{ii}+Mu^{nn}\Big{)}\geq\frac{1}{n}\Big{(}\sum_{i=1}^{n-1}\lambda_{i}^{-1}+M\lambda_{n}^{-1}\Big{)}\geq\Big{(}M\lambda_{1}^{-1}\cdots\lambda_{n}^{-1}\Big{)}^{1/n},\end{split}$
where $\lambda_{1}\leq\cdots\leq\lambda_{n}$ are the eigenvalues of $\{u_{ij}\}$. $c_{3}$ is a positive constant with $c_{3}\leq Cc_{1}$, where $C$ depends on $n$ and the maximum of $u_{t}$. The constant $c_{2}$ depends on the maximum of $u_{t}$, $|Du|$ and the diameter of $\Omega$. Hence, for $A\geq\max\{1,2\underline{a}/c_{1}\}$ and $M\geq((c_{1}+c_{2})/c_{3})^{n}$, we have
(4.9) $\mathcal{L}^{\prime}(A\zeta\pm\mathcal{T}(u-\phi))\geq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times(0,T].$
From now on $M$ is fixed.
Next, we estimate $\zeta$ on $\partial(B_{\varepsilon}\cap\Omega)\times[0,T]$. For $\varepsilon$ small enough, we have
(4.10) $\zeta\leq-\frac{1}{2}\mu\sum_{\beta<n}x_{\beta}^{2}+O(|x^{\prime}|^{3})\leq-\frac{1}{4}\mu|x^{\prime}|^{2},\quad\text{on $(\overline{B_{\varepsilon}}\cap\partial\Omega)\times[0,T]$.}$
Then, on $(\partial B_{\varepsilon}\cap\Omega)\times[0,T]$ and if $\frac{1}{4}\mu|x^{\prime}|^{2}>Mx_{n}^{2}$, we have
(4.11) $\begin{split}\zeta&\leq\frac{1}{2}(B_{\alpha\beta}-\mu\delta_{\alpha\beta})x_{\alpha}x_{\beta}+\frac{1}{2}Mx_{n}^{2}-x_{n}\\ &\leq-\frac{1}{2}\mu|x^{\prime}|^{2}+C|x^{\prime}|^{3}+Mx_{n}^{2}\\ &\leq-c_{4}\varepsilon^{2},\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$
Additionally, on $(\partial B_{\varepsilon}\cap\Omega)\times[0,T]$ and if $\frac{1}{4}\mu|x^{\prime}|^{2}\leq Mx_{n}^{2}$, we have $x_{n}\geq c_{5}\varepsilon$ and
(4.12) $\begin{split}\zeta&\leq C|x^{\prime}|^{2}+Mx_{n}^{2}-x_{n}\\ &\leq(M+C)\varepsilon^{2}-c_{5}\varepsilon\\ &\leq-\frac{1}{2}c_{5}\varepsilon,\quad\quad\text{for $\varepsilon$ small enough.}\end{split}$
Since $|\mathcal{T}(u-\phi)|$ has a upper bound in $\overline{Q}_{T}$, which supposed to be $c_{6}$, then by (4.6), (4.10), (4.11) and (4.12), we get
(4.13) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{on}\quad\partial(B_{\varepsilon}\cap\Omega)\times[0,T],$
for $A=\max\{1,2\underline{a}/c_{1},4\bar{a}/\mu,c_{6}/(c_{4}\varepsilon^{2}),2c_{6}/(c_{5}\varepsilon)\}$. Thus by maximum principle and equations (4.9), (4.7) and (4.13), we get
(4.14) $A\zeta\pm\mathcal{T}(u-\phi)\leq 0,\quad\text{in}\quad(B_{\varepsilon}\cap\Omega)\times[0,T].$
Since $\zeta(0,t_{0})=\mathcal{T}(u-\phi)(0,t_{0})=0$ for each $t_{0}\in[0,T]$, we have
$|\partial_{n}\mathcal{T}(u-\phi)(0,t_{0})|=|u_{\alpha n}(0,t_{0})-\phi_{\alpha n}(0,t_{0})|\leq AN,$
which yields that
(4.15) $|u_{\alpha n}(0,t_{0})|\leq AN+||\phi||_{C^{2}(\overline{Q}_{T})}\leq C.$
By the equation (4.1) and estimate (4.4), we then get $u_{nn}\leq C$ on $\partial\Omega\times[0,T]$ for some universal constant $C$. Combining with the positive lower and upper bounds of $u_{t}$, we have
(4.16) $1/C\leq\lambda(D^{2}u)\leq C\quad\text{on }\partial_{p}Q_{T}$
for some positive constant $C$ depending on $n,\Omega,c_{0},||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and the uniform convexity of $\phi(\cdot,0)$.
(c) global $C^{2}$ estimates. The Legendre transform $U$ of $u$ satisfies
$-U_{t}\det D^{2}U=(1+|y|^{2})^{-\beta}\quad\text{in }Q_{T}^{*}$
with $C^{-1}\leq-U_{t}\leq C$ and $||D^{2}U||_{L^{\infty}(\partial_{p}Q_{T}^{*})}\leq C$. Conducting the same calculation as in Lemma 2.4, we have
$\tilde{L}U_{kk}\geq-2\beta,$
where $\tilde{L}=\frac{1}{U_{t}}\partial_{t}+U^{ij}\partial_{ij}$. Thus $\tilde{L}(U_{kk}-Ct)\geq 0$ for some universal constant $C$. Hence by maximum principle and (4.16),
$\sup_{Q_{T}^{*}}U_{kk}\leq C\Big{(}\sup_{\partial_{p}Q_{T}^{*}}U_{kk}+T\Big{)}\leq C,$
which yields $\lambda(D^{2}u)\geq C$ in $Q_{T}$ for some positive constant $C$ depending on $n,\Omega,c_{0},T,||\phi||_{\tilde{C}^{4}(\overline{Q}_{T})}$ and uniform convexity of $\phi(\cdot,0)$. Therefore we have proved
$1/C\leq\lambda(D^{2}u)\leq C\quad\text{on }\overline{Q}_{T}.$
(d) $\tilde{C}^{2,\alpha}$ estimates. The interior higher regularity of $u$ follows from that of $U$ in $Q_{T}^{*}$ by [26]. Under the same condition of Theorem 2.7, it holds $||u_{t}||_{\tilde{C}^{\alpha}(\overline{Q}_{T})}\leq C$ by Krylov-Safonov’s estimate.
Note that if we let $v=T_{k}(u-\phi)$, $11\leq k\leq n-$ 1, as in subsection 2.3, we have
$-v_{t}+a^{ij}v_{ij}+b^{i}v_{i}=f\quad\text{in }G=B_{R_{0}}^{+}\times(t_{0}-R_{0}^{2},t_{0}].$
The term $b^{i}v_{i}$ comes from the differential of $(1+|Du|^{2})^{-\beta}$ and the related coordinate transformation. Here $a^{ij}$ is positive definite, and $|b^{i}|$, $|v_{i}|$ and $|f|$ are bounded due to the boundedness of $u_{t}$, $Du$ and $D^{2}u$. This is the case we have dealt with in equation (2.22). Thus the $\tilde{C}^{2,\alpha}$ estimate of $u$ on $\overline{\Omega}\times[\epsilon,T]$ follows from the same argument in the proof of Corollary 2.13.
The short-time existence theorem ensures the $\tilde{C}^{2,\alpha}$ estimate of $u$ in a small time interval. The proof in this case is similar to the proof of Theorem 3.1. The Fréchet differential of the map $\Phi$, defined similarly as (3.3), is a linear parabolic operator with extra lower order terms. It is still an isomorphism under some suitable compatibility conditions. Thus the short-time existence of the solution follows from the inverse map theorem.
In conclusion, we have proved the global $\tilde{C}^{2,\alpha}$ estimate for $u$, the solution of equation (4.1). ∎
We also present an existence theorem for equation (4.1).
Theorem 4.2.
Assume $\Omega$ is a smooth and uniformly convex domain in $\mathbb{R}^{n}$, $\phi$ is a smooth function in $\overline{Q}_{T}$ with $\phi_{t}\geq c_{0}>0$ on $\partial\Omega\times[0,T]$, and $\phi(\cdot,0)$ is uniformly convex. If equation (4.1) satisfies the compatibility condition of any order, then there exists a unique smooth solution $u$ to the initial boundary value problem (4.1) that is convex in $x$.
Remark 4.3.
Using the above method, one can also prove the global $\tilde{C}^{2,\alpha}$ estimate for the solutions to the first initial-boundary value problem of $\gamma$-Gauss curvature flow as well as the existence of solutions when $0<\gamma\leq 1$.
Acknowledgement
We would like to thank Professor X.-J. Wang for his suggestions on this topic.